if x = (7-4*3 1/2 ) 1/2 then find the value of (x 3 +1/x 3 ) – 4(x 2 +1/x 4 )+(x+1/x)

if x = (7-4*31/2)1/2  then find the value of (x3+1/x3) – 4(x2+1/x4)+(x+1/x)

Grade:9

2 Answers

Akshay
8 years ago
7-4*rt3 = (2)^2 + (rt3)^2 – 2*rt3*2 = (2-rt3)^2,
x=2-rt3, 1/x=1/(2-rt3)=2+rt3,
x+1/x = 4,
x2 = 7-4*rt3, 1/x2=7+4*rt3,
x2+1/x2=14,
x3=97-56*rt3, //1/x3=97+56*rt3,
x3+1/x3=194,
 
Akshay
8 years ago
sry x3 will ot be 97-56rt3, u can find x3 by multiolying x and x2